OpenStudy (anonymous):

What is the exact area of a circle with a radius of 12 feet?

OpenStudy (anonymous):

@Nurali

OpenStudy (anonymous):

@Nurali

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

english please

OpenStudy (nurali):

Area=pi*r^2 r=12

OpenStudy (anonymous):

not an option

OpenStudy (nurali):

tell me option.

OpenStudy (anonymous):

A. 24π square feet B. 36π square feet C. 121π square feet D. 144π square feet

OpenStudy (nurali):

A=12^2*pi A=144π square feet

OpenStudy (mrnood):

@Nurali it asks for the exact area you should not substitute forpi

OpenStudy (nurali):

Sorry forgot

OpenStudy (mrnood):

(not that pi is 'exact')

OpenStudy (anonymous):

so

OpenStudy (anonymous):

whats the answer

OpenStudy (mrnood):

so you are given the formula pi r^2 r = 12 work it out

OpenStudy (anonymous):

idk no this stuff

OpenStudy (mrnood):

if r is 12 what is r^2 (r squared = r times r) ( = 12 times 12)

OpenStudy (anonymous):

english please

OpenStudy (mrnood):

that IS English what is 12 squared (i.e. 12 times 12)

OpenStudy (anonymous):

144

OpenStudy (anonymous):

and nou were speaking gibrash

OpenStudy (nurali):

Correct

OpenStudy (mrnood):

if oyu think maths notation is "gibrash" then I suggest you concentrate on your English (it's gibberish) You are evidently studying areas and circles If you cannot work out a square of a number then you cannot do this subject. Find out how to do 'squares' and then try to work out pi times 12 squared